Diluting wine.
A restuarant manager has 2 liters of white wine that is 12% alcohol. How many liters of white grape juice should be added to get a drink that is 10 % alcohol?

Answer by ptaylor(2198)

Let x=number of liters of white grape juice that needs to be added
(1) The pure alcohol in the 2 liters of 12% alcohol is (0.12*2).
(2) The pure alcohol in the white grape juice is (0).
(3) The pure alcohol in the final mixture is (0.10(2+x))
Now we know that (1)+(2)=(3). So our equation to solve is:
0.12*2=0.10(2+x) get rid of parens and simplify
0.24=0.20+0.10x subtract 0.20 from both sides
0.24-0.20=0.10x collect like terms
0.04=0.10x divide both sides by 0.10
x=0.4 liter---number of liters of white grape juice that needs to be added
CK
0.12*2=0.10*2.4
0.24=0.24

Hope this helps---ptaylor